If you push a 35 kg box from rest across a floor in 5 seconds and reach a speed of 15m/s, what is the applied force?
1
Expert's answer
2018-05-27T08:43:08-0400
The Newton’s second law states
F=ma
Here the acceleration a is given by
a=∆v/∆t=(15-0)/5=0 m/s^2
So
F=35×3=105 N